# Quantum cryptography beyond key distribution: theory and experiment

### v2

@@ II.1 Conjugate coding
In quantum cryptography with discrete variables, the uncertainty of interest relates to measurements performed in conjugate bases described by the Pauli $X$, $Y$ and $Z$ operators. These describe observables such as photonic polarization, particle spin, spatial and temporal modes, etc. Labeling the eigenstates of the Pauli $Z$ operator as |$\ket{0}$ and |$\ket{1}$, we may express the eigenstates of the $X$ operator as
|$\ket{+}=\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right),$ (2)
|$\ket{-}=\frac{1}{\sqrt{2}}\left(\ket{0}-\ket{1}\right),$
@@
|$\ket{+i}=\frac{1}{\sqrt{2}}\left(\ket{0}+i\ket{1}\right),$ (3)
|$\ket{-i}=\frac{1}{\sqrt{2}}\left(\ket{0}-i\ket{1}\right).$
@@
  • Encode bit 00 as a quantum state uniformly drawn from the set {|$\{\ket{0},\ket{+}\}$.
  • Encode bit $1$ as a quantum state uniformly drawn from the set {|$\{\ket{1},\ket{-}\}$.
@@
Figure 1: Adversarial settings for two-party quantum primitives. a) Honest protocol: Alice generates a quantum state $\rho$ using her source $S$ and an encoding unitary $U_{A}$. Bob applies a unitary $U_{B}$ and performs a measurement with a detection device $D$, possibly followed by some classical post-processing $CPP$   b) Protocol with a dishonest Alice: she can replace $\rho$ with another state $\sigma$, possibly living in a larger Hilbert space   c) Protocol with a dishonest Bob: he can perform any CPTD map on the state $\rho$ he receives.
@@ II.2 No-cloning
Formally, suppose we wish to clone an unknown initial state living in $\mathcal{H}_{i}$ of the form |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$ where $\left(\alpha,\beta\right)\in\mathbb{C}^{2}$ and $|\alpha|^{2}+|\beta|^{2}=1$ onto an initialized copy register living in $\mathcal{H}_{c}$. There exists no unitary $U$ acting on $\mathcal{H}_{i}\otimes\mathcal{H}_{c}$ such that
@@ II.3 Quantum teleportation
The concept of quantum teleportation allows for the transfer of an unknown quantum state |$\ket{\psi}$ using classical communication and pre-shared entanglement Bennett et al. (1993). This concept is present in several cryptographic situations throughout the following sections. For reference, we recall the steps involved:
@@
    The sender makes a Bell-basis measurement of the unknown state |$\ket{\psi}$ together with their half of a singlet.
@@
    The receiver transforms the other half singlet into |$\ket{\psi}$ using one of four unitary operators $I,X,Z,XZ$ as prescribed by $(a,b)$.
Note that among other features, quantum teleportation allows for the reversal of a quantum channel from sender to receiver using a classical channel. If Alice wants to send a qubit to Bob, he can send her a half singlet through a reverse quantum channel, after which Alice can teleport an unknown state |$\ket{\psi}$ in the forward direction using the classical channel.
@@ II.6 Minimal assumptions
Security Assumption Primitive
information -theoretic quantum mechanics (QM) unforgeable tokens, unclonable encryption, data locking, digital signatures, weak coin flipping, cheat-sensitive imperfect mistrustful primitives (Section IV), multipartite primitives (Section VII), private quantum channels, authentication of quantum messages, anonymous communication, blind computation with quantum clients
QM + noisy channel covert communication
QM + special relativity bit commitment, oblivious transfer, strong coin flipping, multiparty computation
QM + bounded/noisy quantum storage data-locked QKD, bit commitment, oblivious transfer, strong coin flipping, multiparty computation
QM + special relativity + bounded entanglement position verification
computational pseudorandom quantum states, one-way state generators, EFI pairs bit commitment, oblivious transfer, strong coin flipping, zero-knowledge proofs
quantum-safe one-way functions encryption with quantum public keys
trapdoor $2$-regular collision-resistant, trapdoor claw-free functions blind computation with classical clients, fully homomorphic encryption
indistinguishability obfuscation public-key money
indistinguishability obfuscation + learning with errors one-shot signatures
hardware-based one-time memories one-time programs
quantum physical unclonable functions device authentication, identification, software licensing
@@ III.1 Unforgeable tokens
Memoryless alternatives were also proposed, achieving spending flexibility through a network of trusted agents and authenticated channels, positioned at precise spatial locations with respect to the verification points Kent and Pitalúa-García (2020); Kent (2019). To some extent, these schemes allow the user to choose the time and place of verification without storing the quantum states. From a security perspective, monitoring the spatial and temporal coordinates of verifiers requires either a trusted Global Positioning System (GPS) or secure position verification schemes discussed in Section III.3. A first partial implementation using weak coherent states was performed in Kent et al. (2022), followed by an experiment demonstrating transaction time advantage over classical schemes Jiang et al. (2024). Note that the work from Kent et al. (2022) also stated three properties that should be simulateneously achieved to claim a quantum advantage: token unforgeability, instant validation and user privacy.
@@ III.3 Position verification (tagging)
The first proposal for authenticating a tagging device’s classical location through the exchange of quantum states was formulated in a patent entitled Tagging systems, whose contents were later published in Kent et al. (2011). In this work, along with independent results from Malaney (2010a, b), several schemes were proposed for i.t.-secure quantum position verification assuming the validity of quantum theory and no-signalling constraints. In FIG. 3, we describe a simple scheme that is secure under the assumption that dishonest provers may not share any entanglement Kent et al. (2011). This assumption is crucial since all previously mentioned schemes are in fact insecure against teleportation-based attacks, in which dishonest provers $P_{0}$ (resp. $P_{1}$) located to the right (resp. left) of $V_{0}$ (resp. $V_{1}$) pre-share a singlet state for each round $i$. As soon as $P_{1}$ receives |$\ket{\psi_{i}}$, they perform a teleportation measurement on their part of the singlet and |$\ket{\psi_{i}}$, forwarding all relevant classical data at light speed to $P_{0}$. As soon as $P_{0}$ receives $b_{i}$, they measure their part of the singlet in basis $b_{i}$, forwarding the measurement basis and outcome to $P_{1}$, who may then retrieve the correct measurement outcome corresponding to the challenged basis Kent et al. (2011); Lau and Lo (2011).
@@
Facing unentangled dishonest provers, the scheme from FIG. 3 cannot tolerate losses beyond $50\%$ / $3$ dB Qi and Siopsis (2015). The reason is similar to that for unforgeable quantum tokens (Section III.1), in which dishonest parties can replace lossy channels by ideal ones: a dishonest $P_{1}$ measures each |$\ket{\psi_{i}}$ in a random basis chosen from $\{Z,X\}$ and waits for dishonest $P_{0}$ to forward the basis bit $b_{i}$. If their random basis choice agrees with $b_{i}$, $P_{1}$ forwards the correct measurement outcome to $V_{1}$, while if their choice disagrees with $b_{i}$, they declare the state as lost. Interestingly, this challenge may be overcome by using more than two encoding bases, uniformly distributed over the Bloch sphere: using a perfect single-photon source and in the limit of many encoding bases, an implementation with $1\%$ error rate may then tolerate up to $14$ dB losses Qi and Siopsis (2015). Other practical schemes were proposed to increase the loss tolerance under various assumptions Allerstorfer et al. (2023, 2022a); Escolà-Farràs and Speelman (2023); Lim et al. (2016); Cowperthwaite et al. (2023).
@@ III.5 Fingerprinting
|$\ket{h_{x}}:=\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\ket{i}\ket{E_{i}(x)},$ (9)
where $E_{i}(x)$ denotes the $i$th bit of $E(x)$. It is easy to see that $\bra{h_{x}}h_{y}\rangle\leq\delta$ where $10<\delta<$ 1 (as given by the distance of codewords in the code) for $x\neq y$, making the $2^{n}$ different quantum states almost orthogonal. The referee in fingerprinting protocols then uses an “equality test” such as the SWAP-test Barenco et al. (1997) to determine whether the two quantum states received are the same (and thus whether the initial bit strings are equal) or different. Since the two states are almost orthogonal, this test succeeds with high probability, and by repeating a constant number of times, the referee fails only with negligible probability to correctly assess two strings’ equality.
@@
Following the proposal from Arrazola and Lütkenhaus (2014), proof-of-principle quantum fingerprinting schemes were experimentally demonstrated, using an altered commercial QKD system at telecom wavelengths Xu et al. (2015) and achieving quantum advantage over $20$km of standard optical fiber Guan et al. (2016). Based on the hidden matching problem introduced in Bar-Yossef et al. (2008), other implementations of quantum advantage for communication complexity were performed, requiring a simple setup consisting of only one beamsplitter and two single-photon detectors Kumar et al. (2019).
@@ III.6 Data locking
A little later, QSDC employing an atomic quantum memory was reported Zhang et al. (2017) and long-distance experiments were provided Zhu et al. (2017); Zhang et al. (2022), along with continuous-variable implementations Cao et al. (2023). For a more complete review of QSDC and its experimental demonstrations, we refer the Reader to Pan et al. (2024).
@@ IV.1 Digital signatures
The first QDS protocol was proposed in Gottesman and Chuang (2001). Based on the classical scheme of Lamport (1979), the main idea was to replace the need for a classical one-way function $f(x)$ (Section II.6) with a quantum state |$\ket{f(x)}$. Lamport’s signatures offer only computational security: with unbounded computational resources, it becomes possible to recover the secret key that contains the pre-image $x$ required to sign a message from the public information $f(x)$ by using brute-force inversion. In contrast, for non-orthogonal quantum states, it is impossible to perfectly recover their classical description (Section II.1). To ensure all security properties listed above (i-iv), the protocol requires a setting-up phase, where parties ensure that they receive the intended quantum states through point-to-point quantum communication and computation. In the signing phase however, each party acts locally, allowing the signing and verification steps to happen without further communication.
@@ IV.2 Bit commitment
Under such computational assumptions, BC comes in two dual flavors: binding but computationally hiding and hiding but computationally binding. Commitments of the first type may be achieved from any one-way function Naor (1991); doi:10.1137/S0097539793244708. The construction of computationally binding and statistically hiding BC from any one-way function was solved more recently in doi:10.1137/080725404.
@@
Despite the no-go result for perfect QBC, i.t.-secure schemes are possible in a setting where both parties can bias the binding and hiding conditions with a non-negligible probability, strictly bounded away from $1$. Some of these schemes have the interesting benefit that any cheating attempt can be detected by the honest parties with a non-zero probability, causing the protocol to abort Hardy and Kent (2004); He (2015); Buhrman et al. (2008); Li et al. (2013). As discussed later in Section IV.4, these protocols are commonly used as a subroutine for cheat-sensitive quantum strong coin flipping Berlín et al. (2009); Pappa et al. (2014); Berlín et al. (2011).
@@ IV.3 Oblivious transfer
In a second form named $1$-out-of-$2$ OT, the sender transfers two messages and the receiver obtains only one message of their choosing Even et al. (1983). In that case, security requires that:
@@
The two versions of OT presented above were proven equivalent in Crépeau (1987). In the context of this review, it is particularly interesting to point out that Stephen Wiesner, who introduced the fundamental idea of conjugate coding in the 1970s (Section II.1), also proposed an early application of quantum cryptography named “multiplexing” which is essentially identical to $1$-out-of-$2$ OT Wiesner (1983). Unfortunately, he failed to see the cryptographic significance of his discovery. While early papers discussed specific cryptographic applications of OT, it is Kilian who first generalized its use to any secure two-party computation Kilian (1988). Noticeable simplifications and efficiency improvements can be found in Crépeau et al. (1995); Ishai et al. (2008).
@@
For completeness, we note the existence of a related primitive known as quantum private queries Giovannetti et al. (2008); Jakobi et al. (2011), in which a user wishes to access an entry from a database while keeping their choice of entry private from the database provider. While secure $1$-out-of-$n$ OT could provide such a service, security can be achieved in a weaker cheat-sensitive model in which the provider does not wish to risk being detected cheating. A fault-tolerant protocol along with its proof-of-principle demonstration were provided in Chan et al. (2014).
@@ IV.4 Strong coin flipping
A naive attempt at performing i.t.-secure SCF with zero bias involves Alice (Bob) generating a private random bit $x_{A}$ ($x_{B}$). At a later pre-agreed time, both parties publicly broadcast their bits and agree to the following flip outcome: $x_{A}\oplus x_{B}$. At first glance, this appears secure since none of the parties can learn or influence the other party’s bit before they communicate. However, simultaneous broadcasting must be enforced, which effectively requires a trusted third party Berlín et al. (2009). Secure classical SCF is nevertheless possible when allowing for a probability $20<\Delta<1/$ 2 that the flip does not provide an outcome Hänggi and Wullschleger (2011), yielding an optimal bias of:
@@
Figure 4: General representation of a quantum coin flipping scheme. Adapted from Aharonov et al. (2016), the protocol starts with a separable state |$\ket{\psi_{0}}$ living in Hilbert space $\mathcal{A}\otimes\mathcal{M}\otimes\mathcal{B}$, where $\mathcal{A}$ and $\mathcal{B}$ are Alice and Bob’s spaces, respectively, and the message register living on $\mathcal{M}$ is sent back and forth between the two parties at each round. At each odd round $i$, Alice applies a unitary $U_{A,i}$ and projection $E_{A,i}$ on the space $\mathcal{A}\otimes\mathcal{M}$, while at every even round $i$, Bob applies a unitary $U_{B,i}$ and projection $E_{B,i}$ on the space $\mathcal{M}\otimes\mathcal{B}$. After the final round $n$, each party measures their private register to obtain the outcome of the coin flip.
Despite these limitations, quantum resources can provide SCF with biases that outperform the best classical bound from Eq. (11). Combining this simple classical benchmark with loss-tolerant quantum SCF proposals Berlín et al. (2009); Pappa et al. (2011), experimental works have demonstrated quantum advantage over optical fiber spools of a few tens of kilometers Pappa et al. (2014); Berlín et al. (2011). These implementations all rely on some form of (imperfect) quantum bit commitment: Alice commits to a bit $x_{A}$ through a single quantum state, while Bob uniformly generates and declares a bit $x_{B}$. During the reveal phase, measurement bases and outcomes are checked for consistency, and the output of the flip is set to $x_{A}\oplus x_{B}$. Although the biases achieved remain high for cryptographic purposes ($\epsilon\approx 0.30-0.40$), these experiments stand as elegant examples of mistrustful quantum schemes outperforming classical schemes despite the established no-go theorems Mayers (1997); Lo and Chau (1997).
@@ IV.6 One-time programs
Using quantum resources, i.t.-secure OTPs with a deterministic output do not exist Broadbent et al. (2013), although allowing for a non-zero error probability in the output circumvents this no-go result Roehsner et al. (2018). In fact, such noisy schemes were demonstrated experimentally, with the four single-bit gates encoded onto the states summarized in Table 2. The input to each program is realized through a projective measurement in the $Z$ basis for 00 and $X$ basis for $1$: this provides a measurement outcome interpreted as the output of the gate, which succeeds with probability $\approx 85.36\%$. The gate rate and loss tolerance of this protocol were later improved using remote state preparation Roehsner et al. (2021).
@@
Constant 00 0     0
1     0 $\frac{1}{\sqrt{2+\sqrt{2}}}\left(\ket{0}+\ket{+}\right)$
Constant $1$ 0     1
1     1 $\frac{1}{\sqrt{2+\sqrt{2}}}\left(\ket{1}-\ket{-}\right)$
@@
1     0 $\frac{1}{\sqrt{2+\sqrt{2}}}\left(\ket{1}+\ket{+}\right)$
Table 2: Summary of the four single bit gates and their corresponding truth tables, along with their quantum encodings proposed in Roehsner et al. (2018). In the middle column, each line indicates a desired input/output relation for the considered gate. Measuring each gate in the $Z$ basis for input 00 and $X$ basis for input $1$ yields a measurement outcome satisfying the truth table with probability $\approx 85.36\%$.

### v3

@@ II.1 Conjugate coding
[…]
@@
[…]
@@
[…]
@@
Figure 1: Adversarial settings for two-party quantum primitives. a) Honest protocol: Alice generates a quantum state $\rho$ using her source $S$ and an encoding unitary $U_{A}$. Bob applies a unitary $U_{B}$ and performs a measurement with a detection device $D$, possibly followed by some classical post-processing CPPCPP   b) Protocol with a dishonest Alice: she can replace $\rho$ with another state $\sigma$, possibly living in a larger Hilbert space   c) Protocol with a dishonest Bob: he can perform any CPTD map on the state $\rho$ he receives.
@@ II.2 No-cloning
[…]
@@ II.3 Quantum teleportation
[…]
@@
[…]
@@
[…]
@@ II.6 Minimal assumptions
Security
Assumption
Primitive
information
-theoretic
quantum mechanics (QM)
unforgeable tokens, unclonable encryption, data locking,
digital signatures, weak coin flipping, cheat-sensitive imperfect
mistrustful primitives (Section IV), multipartite primitives
(Section VII), private quantum channels, authentication of
quantum messages, anonymous communication,
blind computation with quantum clients
QM + noisy channel
covert communication
QM + special relativity
bit commitment, oblivious transfer, strong coin flipping,
multiparty computation
QM + bounded/noisy quantum storage
data-locked QKD, bit commitment, oblivious transfer,
strong coin flipping, multiparty computation
QM + special relativity + bounded entanglement
position verification
computational
pseudorandom quantum states,
one-way state generators,
EFI pairs
bit commitment, oblivious transfer, strong coin flipping,
zero-knowledge proofs
quantum-safe one-way functions
encryption with quantum public keys
trapdoor 22-regular collision-resistant,
trapdoor claw-free functions
blind computation with classical clients,
fully homomorphic encryption
indistinguishability obfuscation
public-key money
indistinguishability obfuscation
+ learning with errors
one-shot signatures
hardware-based
one-time memories
one-time programs
quantum physical unclonable functions
device authentication, identification,
software licensing
@@ III.1 Unforgeable tokens
Memoryless alternatives were also proposed, achieving spending flexibility through a network of trusted agents and authenticated channels, positioned at precise spatial locations with respect to the verification points Kent and Pitalúa-Garc´ıa (2020); Kent (2019). To some extent, these schemes allow the user to choose the time and place of verification without storing the quantum states. From a security perspective, monitoring the spatial and temporal coordinates of verifiers requires either a trusted Global Positioning System (GPS) or secure position verification schemes discussed in Section III.3. A first partial implementation using weak coherent states was performed in Kent et al. (2022), followed by an experiment demonstrating transaction time advantage over classical schemes Jiang et al. (2024). Note that the work from Kent et al. (2022) also stated three properties that should be simulateneously achieved to claim a quantum advantage: token unforgeability, instant validation and user privacy.
@@ III.3 Position verification (tagging)
[…]
@@
[…]
@@ III.5 Fingerprinting
[…]
@@
Following the proposal from Arrazola and Lütkenhaus (2014), proof-of-principle quantum fingerprinting schemes were experimentally demonstrated, using an altered commercial QKD system at telecom wavelengths Xu et al. (2015) and achieving quantum advantage over 2020km of standard optical fiber Guan et al. (2016). Based on the hidden matching problem introduced in Bar-Yossef et al. (2008), other implementations of quantum advantage for communication complexity were performed, requiring a simple setup consisting of only one beamsplitter and two single-photon detectors Kumar et al. (2019).
@@
Finally, we note the existence of other communication complexity tasks that significantly benefit from quantum advantage Raz (1999); Buhrman et al. (1998); Doriguello and Montanaro (2020); Buhrman et al. (2016); Gupta et al. (2023), and refer the Reader to Buhrman et al. (2010) for a more detailed overview.
@@ III.6 Data locking
A little later, QSDC employing an atomic quantum memory was reported Zhang et al. (2017) and long-distance experiments were provided Zhu et al. (2017); Zhang et al. (2022a), along with continuous-variable implementations Cao et al. (2023). For a more complete review of QSDC and its experimental demonstrations, we refer the Reader to Pan et al. (2024).
@@ IV.1 Digital signatures
[…]
@@ IV.2 Bit commitment
Under such computational assumptions, BC comes in two dual flavors: binding but computationally hiding and hiding but computationally binding. Commitments of the first type may be achieved from any one-way function Naor (1991); Håstad et al. (1999). The construction of computationally binding and statistically hiding BC from any one-way function was solved more recently in Haitner et al. (2009).
@@
Despite the no-go result for perfect QBC, i.t.-secure schemes are possible in a setting where both parties can bias the binding and hiding conditions with a non-negligible probability, strictly bounded away from $1$. Some of these schemes have the interesting benefit that any cheating attempt can be detected by the honest parties with a non-zero probability, causing the protocol to abort Hardy and Kent (2004); He (2015); Buhrman et al. (2008); Li et al. (2013). As discussed later in Section IV.4, these protocols are commonly used as a subroutine for cheat-sensitive quantum strong coin flipping Berlín et al. (2009); Pappa et al. (2014); Berl´ın et al. (2011).
@@ IV.3 Oblivious transfer
In a second form named 11-out-of-22 OT, the sender transfers two messages and the receiver obtains only one message of their choosing Even et al. (1983). In that case, security requires that:
@@
The two versions of OT presented above were proven equivalent in Crépeau (1987). In the context of this review, it is particularly interesting to point out that Stephen Wiesner, who introduced the fundamental idea of conjugate coding in the 1970s (Section II.1), also proposed an early application of quantum cryptography named “multiplexing” which is essentially identical to 11-out-of-22 OT Wiesner (1983). Unfortunately, he failed to see the cryptographic significance of his discovery. While early papers discussed specific cryptographic applications of OT, it is Kilian who first generalized its use to any secure two-party computation Kilian (1988). Noticeable simplifications and efficiency improvements can be found in Crépeau et al. (1995); Ishai et al. (2008).
@@
For completeness, we note the existence of a related primitive known as quantum private queries Giovannetti et al. (2008); Jakobi et al. (2011), in which a user wishes to access an entry from a database while keeping their choice of entry private from the database provider. While secure 11-out-of-$n$ OT could provide such a service, security can be achieved in a weaker cheat-sensitive model in which the provider does not wish to risk being detected cheating. A fault-tolerant protocol along with its proof-of-principle demonstration were provided in Chan et al. (2014).
@@ IV.4 Strong coin flipping
A naive attempt at performing i.t.-secure SCF with zero bias involves Alice (Bob) generating a private random bit $x_{A}$ ($x_{B}$). At a later pre-agreed time, both parties publicly broadcast their bits and agree to the following flip outcome: $x_{A}\oplus x_{B}$. At first glance, this appears secure since none of the parties can learn or influence the other party’s bit before they communicate. However, simultaneous broadcasting must be enforced, which effectively requires a trusted third party Berlín et al. (2009). Secure classical SCF is nevertheless possible when allowing for a probability $0<\Delta<1/2$ that the flip does not provide an outcome Hänggi and Wullschleger (2011), yielding an optimal bias of:
@@
[…]
Despite these limitations, quantum resources can provide SCF with biases that outperform the best classical bound from Eq. (11). Combining this simple classical benchmark with loss-tolerant quantum SCF proposals Berlín et al. (2009); Pappa et al. (2011), experimental works have demonstrated quantum advantage over optical fiber spools of a few tens of kilometers Pappa et al. (2014); Berl´ın et al. (2011). These implementations all rely on some form of (imperfect) quantum bit commitment: Alice commits to a bit $x_{A}$ through a single quantum state, while Bob uniformly generates and declares a bit $x_{B}$. During the reveal phase, measurement bases and outcomes are checked for consistency, and the output of the flip is set to $x_{A}\oplus x_{B}$. Although the biases achieved remain high for cryptographic purposes ($\epsilon\approx 0.30-0.40$), these experiments stand as elegant examples of mistrustful quantum schemes outperforming classical schemes despite the established no-go theorems Mayers (1997); Lo and Chau (1997).
@@ IV.6 One-time programs
Using quantum resources, i.t.-secure OTPs with a deterministic output do not exist Broadbent et al. (2013), although allowing for a non-zero error probability in the output circumvents this no-go result Roehsner et al. (2018). In fact, such noisy schemes were demonstrated experimentally, with the four single-bit gates encoded onto the states summarized in Table 2. The input to each program is realized through a projective measurement in the $Z$ basis for 0 and $X$ basis for $1$: this provides a measurement outcome interpreted as the output of the gate, which succeeds with probability $\approx 85.36\%$. The gate rate and loss tolerance of this protocol were later improved using remote state preparation Roehsner et al. (2021).
@@
Constant 0 0     0
[…]
Constant 11 0     1
[…]
@@
[…]
Table 2: Summary of the four single bit gates and their corresponding truth tables, along with their quantum encodings proposed in Roehsner et al. (2018). In the middle column, each line indicates a desired input/output relation for the considered gate. Measuring each gate in the $Z$ basis for input 0 and $X$ basis for input 11 yields a measurement outcome satisfying the truth table with probability $\approx 85.36\%$.